# SceneCraft: Layout-Guided 3D Scene Generation

### v2

@@ 1 Introduction
High-Quality Complex Scene Generation with a 2D Diffusion Model.   Our framework excels in creating 3D scenes by leveraging the advanced generation capabilities of our pretrained 2D diffusion model, SceneCraft2D. SceneCraft2D takes the “bounding-box images (BBI)” rendered from BBS as a condition through ControlNets [76] to generate high-fidelity views of the room that follow the given simple prompt like “This is one view of a [style description] room.” By obtaining high-quality multi-view images through SceneCraft2D, we successfully distill a high-resolution 3D representation [56] of the generated indoor scene.
@@ 2 Related Work
Diffusion-Guided Text-to-3D Generation.   The recent successful 2D diffusion-based generative models [50, 55, 23, 4, 3, 51] have inspired a series of innovative text-to-3D methods [46, 66, 31, 43, 59, 75, 10] to distill powerful 2D pretrained models for 3D content creation. DreamFusion [46] proposes the score distillation sampling (SDS) module to optimize scene representations, e.g., NeRF [41] or Gaussian Splatting [27], of objects by denoising their rendered views. Building on top of DreamFusion, SJC [63], Magic3D [31], and ProlificDreamer [66] alleviate the over-saturation problem and improve the generation quality. Despite impressive results, these methods are restricted in generating small-scale objects without complex semantic composition. More recently, Text2Room [24], SceneScape [17], and Text2NeRF [75] propose to extend object generation to scene generation with off-the-shelf text-image inpainting models [50, 39], where they iteratively inpaint unseen parts of the scene from novel camera perspectives. Another work [37] proposes progressively distilling a text-conditioned indoor panorama generation model [60] using different groups of camera views to optimize a scene representation. However, all of these methods lack semantic control over the generation output other than a simple text prompt. Hence, they cannot be used in the creation of 3D scene models where users want to specify the structure and layouts of the environment. In comparison, our work learns to generate 3D scenes that adhere to user-specified room layouts and textual descriptions, allowing precise control over the environment.
@@ 3 SceneCraft: Methodology
Figure 2: SceneCraft is a novel framework for layout-guided scene generation, which allows users to provide the layout as a bounding-box scene (BBS, Sec. 3.1), a user-friendly layout format that guides the generation. Our framework contains two stages: (a) pretraining of a 2D diffusion model, SceneCraft2D, to solve the 2D version of the layout-guided scene generation task (Sec. 3.2), and (b) distillation of the SceneCraft2D to learn a scene representation of the generated scene (Sec. 3.3).
Our SceneCraft is a novel method for text- and layout-guided scene generation. As illustrated in Figure 2, the input to SceneCraft consists of (1) a prompt as a coarse description of the target scene’s style and content, (2) a “bounding-box scene” (BBS) serving as the layout guidance of the target scene, and (3) a camera trajectory defined in the space of BBS. SceneCraft renders the BBS in the camera trajectory to construct “bounding-box images” (BBI) as the layout condition for a pretrained 2D diffusion model “SceneCraft2D” to generate high-quality 2D images of the scene. With the high-quality images generated by SceneCraft2D, SceneCraft is able to use an SDS-equivalent paradigm [46] to aggregate them into a scene representation (e.g., NeRF [41] or 3D Gaussian splatting [27]) of the generated 3D scene. Notably, our SceneCraft does not require a panoramic view. Instead, our camera view can move freely in the 3D space, enabling the generation of much more complicated indoor layouts consisting of multiple rooms, unlike prior work which only supports single-room scenes.
@@ 3.3 Distillation-Guided Scene Generation
Distillation Process with Annealing.   To generate 3D scenes, we distill the generation ability of our pretrained SceneScraft2D model in a Score Distillation Sampling (SDS) [46, 63]-equivalent pipeline. Unlike the vanilla SDS [46] that works in the latent space and directly works with gradients, our pipeline applies an IN2N [21]-style, which is proven SDS-equivalent by HiFA [78]. In this pipeline, we maintain a multi-view dataset for continual scene representation training while simultaneously and iteratively replacing the multi-view dataset with newly generated images by SceneCraft2D. Through this process, the multi-view dataset will be gradually replaced with views of the generated scenes, which are used to fit the scene representation towards the generation.
@@ 4 Experiments
Implementation and Datasets.   For the development of our SceneCraft2D diffusion model, we finetune Stable Diffusion [50] with our produced layout data. We use multi-view images from ScanNet++ [72] and HyperSim [49] to construct BBI data. In the distillation process, we choose Nerfacto from NeRFStudio [56] as our backbone for scene representation. During distillation, we use a duo-GPU pipeline to parallelize diffusion generation and NeRF training. More details are provided in Appendix Sec. A.
@@ 4.2 Ablation Study
Figure 6: Effect of Base Prompt. Using our base prompt successfully avoids the overfitting and maintains the inherent power of pretrained Stable Diffusion, while using BLIP2 captions leads to control failure.
@@ 5 Conclusion
Figure A: An illustration of duo-GPU training scheduling.
@@ Appendix A Additional Details
Data Processing.   As mentioned in Sec. 4, we use two datasets: HyperSim [49] and ScanNet++ [72], which are synthetic data and real-world data, respectively. We mainly use the HyperSim dataset in our experiments. For HyperSim, the original data provide approximately 77400 images of 461 indoor scenes, with the corresponding camera parameters and semantic bounding boxes. However, some of them do not meet the quality requirements of our training. Following prior work [53], we examine the entire dataset and filter out the lower-quality portions, such as rooms containing extremely complex and irregularly shaped objects, unbounded outdoor space, and rooms with excessively large scales (e.g., churches, restaurants), ultimately retaining approximately half of the original data, amounting to around 24k pairs. ScanNet++ shows more complicated scenes (450+ scenes) compared with HyperSim. We voxelize them with a unit size equal to $0.2m$ to make the trade-off between rendering cost and data quality. Our processed data are publicly available  [^1] Layout Scannet++: https://huggingface.co/datasets/gzzyyxy/layout_diffusion_scannetpp_voxel0.2  [^2] Layout Hypersim: https://huggingface.co/datasets/gzzyyxy/layout_diffusion_hypersim. We leverage the Ray-OBB model (extended from the Ray-AABB model [26]) to complete the rasterization process.
Duo-GPU Training Scheduling.   Inspired by [8, 42], we have implemented a duo-GPU scheduler (see Figure A) for efficient diffusion and NeRF generation. Specifically, the second (or any GPU other than the first) GPU continuously generates new images to update the dataset, while the first GPU continuously trains Nerfacto with the current dataset. Once the diffusion procedure needs images from the first GPU to refine, the first GPU will switch to an offline renderer. This configuration effectively decouples the diffusion generation process, which is inherently more time-intensive, from the comparatively rapid NeRF training, thus streamlining the overall distillation workflow without compromising on quality or efficiency.
@@ B.1 Ablation Study
Effect of Layout-Aware Depth Constraint.   We conduct experiments to validate the effectiveness of the layout-aware depth constraint. As depicted in Figure B, without this strategy, the model is completely unable to learn the correct geometry of the scene from 2D guidance. Although a reasonable appearance is achieved, it is mainly due to the fully flexible camera trajectories and complex layout conditions. Notably, using fixed camera poses as in panorama generation can partially alleviate this issue. With the layout-aware depth constraint, the scene geometry rapidly converges to the ground truth during the very initial training stage. Although some area may find incorrect positions (red box), they are corrected during subsequent training after intermediately disabling this strategy, eventually leading to final convergence (green box). Once the geometry converges well, the color and texture begin to emerge.
@@ Appendix E License of Dataset Used
In this section, we list the licenses of all the datasets we have used during our evaluation:
In addition, we utilize a number of public foundation model checkpoints pretrained on various data sources in our paper. Please refer to their original papers [50, 39] for the license of datasets they have used in pretraining their models.

### v3

@@ 1 Introduction
High-Quality Complex Scene Generation with a 2D Diffusion Model.   Our framework excels in creating 3D scenes by leveraging the advanced generation capabilities of our pre-trained 2D diffusion model, SceneCraft2D. SceneCraft2D takes the “bounding-box images (BBI)” rendered from BBS as a condition through ControlNets [76] to generate high-fidelity views of the room that follow the given simple prompt like “This is one view of a [style description] room.” By obtaining high-quality multi-view images through SceneCraft2D, we successfully distill a high-resolution 3D representation [56] of the generated indoor scene.
@@ 2 Related Work
Diffusion-Guided Text-to-3D Generation.   The recent successful 2D diffusion-based generative models [50, 55, 23, 4, 3, 51] have inspired a series of innovative text-to-3D methods [46, 66, 31, 43, 59, 75, 10] to distill powerful 2D pre-trained models for 3D content creation. DreamFusion [46] proposes the score distillation sampling (SDS) module to optimize scene representations, e.g., NeRF [41] or Gaussian Splatting [27], of objects by denoising their rendered views. Building on top of DreamFusion, SJC [63], Magic3D [31], and ProlificDreamer [66] alleviate the over-saturation problem and improve the generation quality. Despite impressive results, these methods are restricted in generating small-scale objects without complex semantic composition. More recently, Text2Room [24], SceneScape [17], and Text2NeRF [75] propose to extend object generation to scene generation with off-the-shelf text-image inpainting models [50, 39], where they iteratively inpaint unseen parts of the scene from novel camera perspectives. Another work [37] proposes progressively distilling a text-conditioned indoor panorama generation model [60] using different groups of camera views to optimize a scene representation. However, all of these methods lack semantic control over the generation output other than a simple text prompt. Hence, they cannot be used in the creation of 3D scene models where users want to specify the structure and layouts of the environment. In comparison, our work learns to generate 3D scenes that adhere to user-specified room layouts and textual descriptions, allowing precise control over the environment.
@@ 3 SceneCraft: Methodology
Figure 2: SceneCraft is a novel framework for layout-guided scene generation, which allows users to provide the layout as a bounding-box scene (BBS, Sec. 3.1), a user-friendly layout format that guides the generation. Our framework contains two stages: (a) pre-training of a 2D diffusion model, SceneCraft2D, to solve the 2D version of the layout-guided scene generation task (Sec. 3.2), and (b) distillation of the SceneCraft2D to learn a scene representation of the generated scene (Sec. 3.3).
Our SceneCraft is a novel method for text- and layout-guided scene generation. As illustrated in Figure 2, the input to SceneCraft consists of (1) a prompt as a coarse description of the target scene’s style and content, (2) a “bounding-box scene” (BBS) serving as the layout guidance of the target scene, and (3) a camera trajectory defined in the space of BBS. SceneCraft renders the BBS in the camera trajectory to construct “bounding-box images” (BBI) as the layout condition for a pre-trained 2D diffusion model “SceneCraft2D” to generate high-quality 2D images of the scene. With the high-quality images generated by SceneCraft2D, SceneCraft is able to use an SDS-equivalent paradigm [46] to aggregate them into a scene representation (e.g., NeRF [41] or 3D Gaussian splatting [27]) of the generated 3D scene. Notably, our SceneCraft does not require a panoramic view. Instead, our camera view can move freely in the 3D space, enabling the generation of much more complicated indoor layouts consisting of multiple rooms, unlike prior work which only supports single-room scenes.
@@ 3.3 Distillation-Guided Scene Generation
Distillation Process with Annealing.   To generate 3D scenes, we distill the generation ability of our pre-trained SceneScraft2D model in a Score Distillation Sampling (SDS) [46, 63]-equivalent pipeline. Unlike the vanilla SDS [46] that works in the latent space and directly works with gradients, our pipeline applies an IN2N [21]-style, which is proven SDS-equivalent by HiFA [78]. In this pipeline, we maintain a multi-view dataset for continual scene representation training while simultaneously and iteratively replacing the multi-view dataset with newly generated images by SceneCraft2D. Through this process, the multi-view dataset will be gradually replaced with views of the generated scenes, which are used to fit the scene representation towards the generation.
@@ 4 Experiments
Implementation and Datasets.   For the development of our SceneCraft2D diffusion model, we finetune Stable Diffusion [50] with our produced layout data. We use multi-view images from ScanNet++ [72] and HyperSim [49] to construct BBI data. In the distillation process, we choose Nerfacto from NeRFStudio [56] as our backbone for scene representation. During distillation, we use a dual-GPU pipeline to parallelize diffusion generation and NeRF training. More details are provided in Appendix Sec. A.
@@ 4.2 Ablation Study
Figure 6: Effect of Base Prompt. Using our base prompt successfully avoids the overfitting and maintains the inherent power of pre-trained Stable Diffusion, while using BLIP2 captions leads to control failure.
@@ 5 Conclusion
Acknowledgments
This work was supported in part by NSF Grant 2106825, NIFA Award 2020-67021-32799, the IBM-Illinois Discovery Accelerator Institute, the Toyota Research Institute, and the Jump ARCHES endowment through the Health Care Engineering Systems Center at Illinois and the OSF Foundation. This work used computational resources on NCSA Delta through allocations CIS220014 and CIS230012 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, and on TACC Frontera through the National Artificial Intelligence Research Resource (NAIRR) Pilot.
@@ Acknowledgments
Figure A: An illustration of dual-GPU training scheduling.
@@ Appendix A Additional Details
Data Processing.   As mentioned in Sec. 4, we use two datasets: HyperSim [49] and ScanNet++ [72], which are synthetic data and real-world data, respectively. We mainly use the HyperSim dataset in our experiments. For HyperSim, the original data provide approximately 77400 images of 461 indoor scenes, with the corresponding camera parameters and semantic bounding boxes. However, some of them do not meet the quality requirements of our training. Following prior work [53], we examine the entire dataset and filter out the lower-quality portions, such as rooms containing extremely complex and irregularly shaped objects, unbounded outdoor space, and rooms with excessively large scales (e.g., churches, restaurants), ultimately retaining approximately half of the original data, amounting to around 24k pairs. ScanNet++ shows more complicated scenes (450+ scenes) compared to HyperSim. We voxelize them with a unit size equal to $0.2m$ to make the trade-off between rendering cost and data quality. Our processed data are publicly available  [^1] Layout Scannet++: https://huggingface.co/datasets/gzzyyxy/layout_diffusion_scannetpp_voxel0.2  [^2] Layout Hypersim: https://huggingface.co/datasets/gzzyyxy/layout_diffusion_hypersim. We leverage the Ray-OBB model (extended from the Ray-AABB model [26]) to complete the rasterization process.
Dual-GPU Training Scheduling.   Inspired by [8, 42], we have implemented a dual-GPU scheduler (see Figure A) for efficient diffusion and NeRF generation. Specifically, the second (or any GPU other than the first) GPU continuously generates new images to update the dataset, while the first GPU continuously trains Nerfacto with the current dataset. Once the diffusion procedure needs images from the first GPU to refine, the first GPU will switch to an offline renderer. This configuration effectively decouples the diffusion generation process, which is inherently more time-intensive, from the comparatively rapid NeRF training, thus streamlining the overall distillation workflow without compromising on quality or efficiency.
@@ B.1 Ablation Study
Effect of Layout-Aware Depth Constraint.   We conduct experiments to validate the effectiveness of the layout-aware depth constraint. As depicted in Figure B, without this strategy, the model is completely unable to learn the correct geometry of the scene from 2D guidance. Although a reasonable appearance is achieved, it is mainly due to the fully flexible camera trajectories and complex layout conditions. Notably, using fixed camera poses as in panorama generation can partially alleviate this issue. With the layout-aware depth constraint, the scene geometry rapidly converges to the ground truth during the very initial training stage. Although some areas may find incorrect positions (red box), they are corrected during subsequent training after intermediately disabling this strategy, eventually leading to final convergence (green box). Once the geometry converges well, the color and texture begin to emerge.
@@ Appendix E License of Dataset Used
In this section, we list the licenses of all the datasets we have used during our evaluation.
In addition, we utilize a number of public foundation model checkpoints pre-trained on various data sources in our paper. Please refer to their original papers [50, 39] for the license of datasets they have used in pre-training their models.